# Dynamical stability for dense patterns in discrete attractor neural networks

### v1

@@ Abstract
Neural networks storing multiple discrete attractors are canonical models of biological memory. Previously, the dynamical stability of such networks could only be guaranteed under highly restrictive conditions. Here, we derive a theory of the local stability of discrete fixed points in a broad class of networks with graded neural activities and in the presence of noise. By directly analyzing the bulk and outliers of the Jacobian spectrum, we show that all fixed points are stable below a critical load that is distinct from the classical critical capacity and depends on the statistics of neural activities in the fixed points as well as the single-neuron activation function. Our analysis highlights the computational benefits of threshold-linear activation and sparse-like patterns.
@@ Introduction.
In this letter, we extend the ‘Gardner’ approach to gain analytical insights about the stability of fixed points in a broad class of networks with graded neural activities and generic, non-saturating, rectified, power-law activation functions. In particular, rather than the oft-studied sparse limit, here we consider dense patterns for analytical tractability on dynamical stability, and also because the inherent noisiness of neural signaling can easily prevent firing rates from being exactly zero in practice. To supplement these analyses, we also consider sparse patterns and show that our results extend to those in numerical simulations, and in some cases even analytically. We demonstrate that there is a phase transition for stability in such networks: optimizing network connectivity to maintain memory patterns as fixed points with minimal weights renders either all or none of those fixed points stable, depending on pattern statistics and single-neuron properties. We thus characterize the conditions under which fixed-point stability emerges in a large class of network dynamics, providing design principles for biological systems performing auto-associative memory [14, 15].
@@
Figure 1: Storing fixed-points. (a) Color code for combinations of I/O function exponent and smoothness (log-scale). The activation function for selected combinations is illustrated (insets, top). These combinations are used in (b)-(d). (b,c) The sum (b) and L2-norm (c) of the presynaptic weights of a neuron (i.e. a row of $\mathbf{W}$) as a function of memory load ($\alpha$) for different activation functions (colors as in (a)), and $\textrm{CV}=1$. The inset shows divergence of L2-norm as $\alpha\rightarrow 1$, as predicted by theory (note linear scale for the y-axis and extended range on the x-axis). (d) The fraction of patterns which are correctly stored (thick curves) or stable for recall (thin curves) as a function of memory load ($\alpha$) for different activation functions (colors as in a-c), and $\textrm{CV}=2$. (e) Probability density of log-normal distributed patterns at different values of CV (color coded). The inset shows normalized density on a log scale. (f-h) Same as b-d for different values of pattern CV (colors as in e) and $\sigma=1$, $n=1.5$ (f,g); or $\sigma=1$, $n=1$ (h). Solid vs. dotted lines in (b)-(d), (f)-(h) show numerical vs. theoretical results. Numerical simulations optimized the weights $\mathbf{W}$ according to Eq. 4, and also optimized the threshold $\theta$ using finite differentiation. Note that only numerical results are shown for stability (d) and (h) – see subsequent figures for corresponding theoretical results.
Figure 2: Fixed-point stability. (a) Example spectrum - overlaid eigenvalues of $P$ Jacobians $\mathbf{J_{\mu}}$ of a single problem (parameters indicated by text). Annotations of the theoretical values $\lambda_{\mathrm{bulk}}$ (blue line), $\lambda_{\mathrm{mem}}$ (orange line), and $\lambda_{\mathrm{ave}}$ (green line). Inset shows a near-linear relation between $r^{\mu}_{i}$ (x-axis) and $f_{0}(r^{\mu}_{i})$ (y-axis), a prerequisite for an outlier $\lambda_{\mathrm{mem}}$. (b) Same as (a), for six parameter sets; columns differ by the maximal theoretical value. (c-e) Probability densities at different CV values (color-coded) - the difference between the empirical and predicted value of $\lambda_{\mathrm{bulk}}$ (c); the difference between the empirical spectral abscissa $\lambda^{*}$ (the real part of the eigenvalue with the largest real part) and the predicted value of $\lambda_{\mathrm{ave}}$ when varying the threshold $\theta$ such that $\lambda_{\mathrm{ave}}$ becomes the spectral abscissa (d); the difference between the empirical and predicted value of $\lambda_{\mathrm{mem}}$ (e).
@@
We study a network model of $N$ neurons with voltage dynamics:
@@
where $v_{i}$ is the voltage of neuron $i$, $\tau$ is the neural time constant (assumed to be shared across neurons), $\mathbf{W}\in\mathbb{R}^{N\times N}$ defines recurrent connection weights (such that $W_{ij}$ is the strength of the connection from neuron $j$ to neuron $i$), $g_{i}\!\left(\mathbf{v}\right)=g\!\left(v_{i}\right):\mathbb{R}\rightarrow\mathbb{R}^{+}$ is the neural activation function (also shared across neurons) that maps the ‘voltage’ of a neuron, $v_{i}$, to its (positive) instantaneous firing rate, and $\theta_{i}=\theta\in\mathbb{R}$ is the ‘threshold’ (or negative bias) for neuron $i$ (also shared across neurons). Defining $\mathbf{v}=\mathbf{W}\,\mathbf{r}$, Eq. 1 has an equivalent form of rate dynamics [16]
@@
which is the form we will use in the following for mathematical convenience.
@@
$\mathbf{W}^{*}=\mathop{\textrm{argmin}}_{\mathbf{W}}\left\|\mathbf{W}\right\|_{\mathrm{F}}\text{ s.t. }\mathbf{r}^{\mu}=\mathbf{g}\left(\mathbf{W}\,\mathbf{r}^{\mu}\right)\;\forall\mu=1\ldots P$ (4)
We consider the additional constraint that there are no self-couplings, i.e. the diagonal elements of the connection weight matrix are zero, $\mathrm{diag}\left(\mathbf{W}\right)=\boldsymbol{0}$.
We deviate from previous approaches [1, 2, 5, 12] in two important ways. First, we not only require memory patterns to be merely fixed points of the dynamics (which will be guaranteed once Eq. 4 is solved), but we also study the local stability of these fixed points – this is critical for a well-functioning auto-associative memory if it is to perform memory recall by pattern completion (converging to a memory pattern when started from a state that is near but not identical to it) [1, 17]. Specifically, we analyse the Jacobian of the dynamics around each memory pattern
@@
[…]
To summarise, we are interested in how the maximal number of patterns $P$ that can be stored as fixed points, and the fraction of these fixed points that are dynamically stable, depends on parameters $\sigma$, $n$, $\theta$, and CV (as well as $f$, the sparseness of patterns, Details). For obtaining this maximum, we optimize $\mathbf{W}$ per Eq. 4.
@@
A replica analysis of the solution for the optimization problem Eq. 4 follows the Gardner approach [11, 12, 18], without assuming specific connection weights as in the Hopfield model [2, 5]. In the Appendix, we derive a mean-field theory for the capacity to store graded, random memory patterns as fixed points of Eq. 1, without assumptions on the weights. The analysis applies to either dense patterns with a strictly monotonic activation function or to sparse patterns with a rectified monotonic activation function. It generalizes previous work in which a rectified-linear activation function and normalized connection weights were assumed [12]. As usual, our analysis assumes $P,N\to\infty$ with a fixed memory load $\alpha=P/N$ and uses the replica technique and assuming the replica symmetry ansatz.
The resulting theory predicts that the critical capacity $\alpha_{\mathrm{C}}$ depends only on pattern sparseness and is independent of pattern statistics (CV; Fig. 1e) and activation function details (smoothness $\sigma$ and exponent $n$; Eq. 3, Fig. 1a), thus extending results from the rectified-linear case (corresponding to $\sigma=0$ and $n=1$) to our broader class of activation functions ($\sigma\geq 0$ and $n>0$) [12]. Results for the sparse case are presented in Fig. 5 (Details). In the dense case, the critical capacity is $\alpha_{\mathrm{C}}=1$ (note its independence from any of the parameters), and theory predicts the following moments for each element of the connection weight matrix:
@@
For a subcritical load, $\alpha=P/N<\alpha_{\mathrm{C}}$, theory can be compared with numerical experiments by directly solving Eq. 4 using off-the-shelf optimizers [19]. The theory provides a good match for both the mean ($N\,\left\langle W\right\rangle$ from Eq. 7; Fig. 1b,f and Fig. 5c) and the L2-norm of a row of the connection weight matrix ($\sqrt{N\,\left\langle W^{2}\right\rangle}$ from Eq. 6; Fig. 1c,g, and Fig. 5d). As $\alpha$ approaches $\alpha_{\mathrm{C}}$, there is no longer a solution to Eq. 4 (Fig. 1d,h and Fig. 5e), and the L2-norm of weights diverges, as predicted by theory (Fig. 1c,g, inset). Importantly, investigating the emergent stability of patterns in the subcritical regime reveals a second phase transition: all patterns tend to be stable up to a smaller critical value $\alpha_{\mathrm{S}}$, above which all patterns tend to be unstable (Fig. 1d,h and Fig. 6e,f). In the following, we investigate this phenomenon more closely.
@@
where $\mathbf{R},\mathbf{V}\in\mathbb{\mathbb{R}}^{N\times P}$ are the stored memory patterns $R_{i\mu}=r_{i}^{\mu}$ and $V_{i\mu}=g^{-1}(R_{i\mu})+\theta$. A slightly more complicated expression is available when avoiding self-couplings, Eq. 18, but we find that we can neglect this difference and use Eq. 8 in the derivation of stability. This generalizes earlier results derived for storing binary patterns, in which case weights with a similar structure, defined through the patterns’ pseudo-inverse, were already proposed in the ‘Hebbian’ approach[20, 3], but this was limited to a linear activation function $\mathbf{V}=\mathbf{R}$, rendering $\mathbf{W}$ symmetric (and thus allowing the construction of an energy function).
To characterize the critical load for stability, $\alpha_{\mathrm{S}}$, we previously analyzed the eigenvalue spectrum of random matrices $\mathbf{X}\,\mathbf{Y}^{\dagger}$ when pairs of the corresponding entries of $\mathbf{X}$ and $\mathbf{Y}$ are (jointly) independent and identically distributed (i.i.d.), correlated Gaussian [21]. This provides an exact result for the support of the eigenvalue spectrum (Fig. 7), and expressions for the largest real eigenvalue $\lambda_{\mathrm{bulk}}$ (Eq. 20), which we use to derive the critical load for stability, Eq. 21 (Details). For the Jacobian (Eq. 5) this is only an approximation, since the entries of the Jacobian are neither i.i.d. (due to the $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r^{\mu}}\right)\right)$ term which couples different rows), nor Gaussian (since the first term is a deterministic function of the second term, rather than being a jointly distributed random variable). Nevertheless, our empirical results below suggest that it is an acceptable approximation at least of the bulk of the eigenvalue spectrum, with some notable outliers that we will analyze separately below. The resulting critical load for stability from this analysis, describing the bulk of the eigenvalue spectrum, is:
@@
expressed in terms of three cross-correlation terms $c_{\mathrm{rf}}=\left\langle\delta r\,\delta f\!\left(r\right)\right\rangle$, $c_{\mathrm{ff}}=\langle\delta f\!\left(r\right)^{2}\rangle$, and $c_{\mathrm{rr}}=\langle\delta r^{2}\rangle$, where $f\!\left(r\right)=\langle g^{\prime}\!\left(g^{-1}\!\left(r^{\prime}\right)\right)\rangle_{r^{\prime}}\left(g^{-1}\!\left(r\right)+\theta\right)$, with the averaging with respect to an independent variable $r^{\prime}$ neglecting the statistical dependence between different rows (i.e., ignoring the pattern-specific part of the Jacobian).
We note that $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ depends on $\theta$ only through $c_{\mathrm{ff}}$ and is maximized when $c_{\mathrm{ff}}$ is minimized, which occurs at:
$\theta^{*}=-\langle g^{-1}\!\left(r\right)\rangle$ (10)
so by Eq. 6 the average weight becomes 00 for this choice. Furthermore, we note that as long as the smoothness $\sigma$ is not too high, the threshold is expected to be negative.
@@
The foregoing stability analysis was based on the zero-crossing of the ‘rightmost point’ of the bulk of the eigenvalue spectrum, $\lambda_{\mathrm{bulk}}$ (Fig. 2a). However, we empirically find that while $\lambda_{\mathrm{bulk}}$ often predicts stability well (Fig. 2c and Fig. 2b, left column), there are also cases when it alone is an imperfect predictor of stability (Fig. 2b, middle and right columns). Thus, we analyze two specific outlier eigenvalues that might interfere with stability.
The first outlier eigenvalue is associated with the uniform vector as a ‘counterfactual’ eigenvector. If such an eigenvector existed, the corresponding eigenvalue would be the (average) row-sum of the Jacobian, $\bar{J}$. The expected value of this outlier eigenvalue is thus:
$\lambda_{\mathrm{ave}}=\left\langle\bar{J}\right\rangle=-1+\left\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\right\rangle\,N\,\left\langle W\right\rangle$ (11)
where, by taking the expectation, we have ignored (again) the memory pattern-dependence of $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)$ and $N\,\langle W\rangle$ can be computed using replica theory (Eq. 7):
@@
$\lambda_{\mathrm{ave}}=-1+\left\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\right\rangle\,\frac{\theta+\left\langle g^{-1}\!\left(r\right)\right\rangle}{\langle r\rangle}$ (12)
Even though the uniform vector is not actually an eigenvector of the Jacobian in general, we empirically find that eigenvalue(s) often still exist close to $\lambda_{\mathrm{ave}}$ (Fig. 2a), and in fact these can even determine the dynamical stability of the Jacobian (Fig. 2d and Fig. 2b, middle column). Finally, requiring $\lambda_{\mathrm{ave}}<0$, Eq. 12 yields a constraint on the threshold $\theta$:
$\theta<-\langle g^{-1}\!\left(r\right)\rangle+\langle r\rangle/\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\rangle$ (13)
which is always satisfied using $\theta^{*}$ from Eq. 10.
The second Jacobian matrix outlier is associated with the memory pattern itself. If for some constant $c$, $\mathbf{r^{\mu}}+c\mathbf{1}$ is an eigenvector of the Jacobian $\mathbf{J_{\mu}}$, it must satisfy:
$\mathbf{f_{0}\left(r^{\mu}\right)}=\left(\lambda_{\mathrm{mem}}^{\mu}+1\right)\mathbf{r^{\mu}}+c\left(\lambda_{\mathrm{mem}}^{\mu}-\lambda_{\mathrm{ave}}\right)\mathbf{1}$ (14)
where $f_{0}\left(x\right)=g^{\prime}\!\left(g^{-1}\!\left(x\right)\right)\left(g^{-1}\!\left(x\right)+\theta\right)$, and the linear relation between $\mathbf{r^{\mu}}$ and $\mathbf{f_{0}\left(r^{\mu}\right)}$ is captured by (Details):
$\tau_{\mathrm{mem}}^{\mu}$ $=\mathrm{corrcoef}\left(\mathbf{r^{\mu}},\mathbf{f_{0}\left(r^{\mu}\right)}\right)\approx 1$ (15)
$\lambda_{\mathrm{mem}}^{\mu}$ $=-1+c_{rf0}^{\mu}/c_{rr}$ (16)

### v3

@@ Abstract
Neural networks storing multiple discrete attractors are canonical models of biological memory. Previously, the dynamical stability of such networks could only be guaranteed under highly restrictive conditions. Here, we derive a theory of the local stability of discrete fixed points in a broad class of networks with graded neural activities and in the presence of noise. By directly analyzing the bulk and the outliers of the Jacobian spectrum, we show that all fixed points are stable below a critical load that is distinct from the classical critical capacity and depends on the statistics of neural activities in the fixed points as well as the single-neuron activation function. Our analysis highlights the computational benefits of threshold-linear activation and sparse-like patterns.
@@ Introduction.
In this letter, we extend the ‘Gardner’ approach to gain analytical insights about the stability of fixed points in a broad class of networks with graded neural activities and generic, non-saturating, rectified, power-law activation functions. In particular, rather than the oft-studied sparse limit, here we consider dense patterns for analytical tractability on dynamical stability, and also because the inherent noisiness of neural signaling can easily prevent firing rates from being exactly zero in practice. To supplement these analyses, we also consider sparse patterns (Fig. S1a) and show that our results extend to those in numerical simulations. We demonstrate that there is a phase transition for stability in such networks: optimizing network connectivity to maintain memory patterns as fixed points with minimal weights renders either all or none of those fixed points stable, depending on pattern statistics and single-neuron properties. By characterizing the conditions under which fixed-point stability emerges in a large class of network dynamics, we provide design principles for biological systems performing auto-associative memory [14, 15].
@@
Figure 1: Storing fixed-points. (a) Color code for combinations of I/O exponent (x-axis) and smoothness (y-axis, log-scale). The activation function for selected combinations is illustrated (insets, top). These combinations are used in b-e, with $\textrm{CV}=2,\theta=-2$. (b,c) The sum (b) and L2-norm (c) of the presynaptic weights of a neuron (i.e. a row of $\mathbf{W}$) as a function of memory load ($\alpha$) for different activation functions (colors as in a). Note the divergence of L2-norm as $\alpha\rightarrow 1$, as predicted by theory. (d) Measures of the weights asymmetry (top) and dynamics non-normality (bottom) at different load values (x-axis), for different activation functions (colors as in a). (e) The fraction of patterns which are correctly stored (thick curves) or stable for recall (thin curves) as a function of memory load ($\alpha$) for different activation functions (colors as in a). (f) Probability density of log-normal distributed patterns at different values of CV (color coded). The inset shows normalized density on a log scale. (g-j) Same as b-e for different values of pattern CV (colors as in f) with $\sigma=1,n=1,\theta=-2$. Solid vs. dotted lines in b,c,e, and g,h,i show numerical vs. theoretical results (with only numerical results shown for stability e,j – see subsequent figures for corresponding theoretical results). Numerical simulations optimized the weights $\mathbf{W}$ according to Eq. 4 at $N=256$.
@@
We study a network of $N$ neurons with voltage dynamics:
@@
where $v_{i}$ is the voltage of neuron $i$, $\tau$ is the neural time constant, $\mathbf{W}\in\mathbb{R}^{N\times N}$ defines recurrent connection weights (such that $W_{ij}$ is the strength of the connection from neuron $j$ to neuron $i$), $g_{i}\!\left(\mathbf{v}\right)=g\!\left(v_{i}\right):\mathbb{R}\rightarrow\mathbb{R}^{+}$ is the neural activation function that maps the ‘voltage’ of a neuron, $v_{i}$, to its (positive) instantaneous firing rate, and $\theta_{i}=\theta\in\mathbb{R}$ is the ‘threshold’ (or negative bias) for neuron $i$. (Note that $\tau$, $g\!\left(\cdot\right)$, and $\theta$ are shared across neurons.) Defining $\mathbf{v}=\mathbf{W}\,\mathbf{r}$, Eq. 1 has an equivalent form of rate dynamics [16], which is the form we will use in the following for mathematical convenience:
@@
$\begin{array}[]{c}\displaystyle\mathbf{W}^{*}=\mathop{\textrm{argmin}}_{\mathbf{W}}\left\|\mathbf{W}\right\|_{\mathrm{F}}\\[10.0pt] \displaystyle\text{ s.t. }\mathbf{r}^{\mu}=\mathbf{g}\!\left(\mathbf{W}\,\mathbf{r}^{\mu}-\boldsymbol{\theta}\right)\;\forall\mu=1\ldots P\end{array}$ (4)
We consider the additional constraint that there are no self-couplings, i.e. the diagonal elements of the connection weight matrix are zero.
We deviate from previous approaches [1, 2, 5, 12] in two important ways. First, we not only require memory patterns to be merely fixed points of the dynamics (which will be guaranteed once Eq. 4 is solved), but we also study the local stability of these fixed points – this is critical for a well-functioning auto-associative memory if it is to perform memory recall by converging to a memory pattern when started from a state that is near but not identical to it [1, 17]. Specifically, we analyze the Jacobian of the dynamics around each memory pattern
@@
Second, unlike previous approaches that considered binary [1, 2] or sparse memory patterns [5, 12], here we specifically focus on dense patterns in which firing rates are never exactly zero: $\mathbf{r}^{\mu}\in\mathbb{\mathbf{R}}_{+}^{N}$. Again, to allow a systematic study of how stability depends on the properties of this distribution, we consider here the specific case of a log-normal distribution of patterns, which we parametrize by its coefficient of variation $\textrm{CV}=\sqrt{\left\langle\delta r^{2}\right\rangle}/\left\langle r\right\rangle$), while fixing its mean at $\left\langle r\right\rangle=1$ (Fig. 1f), where here and in the following $\left\langle\cdot\right\rangle$ and $\delta$ denote averaging across the distribution of memory patterns and computing the deviation from such an average, respectively. For a hard-rectified power law activation function (i.e. $\sigma=0$, Fig. S1b), scaling the mean of memory patterns while keeping their CV constant would cancel in the Jacobian of the dynamics, Eq. 5, thus leaving stability unaffected. For our soft-rectified activation function, this does not hold exactly, but we find the effects on stability to be negligible, Fig. S4. While such patterns are always technically dense, they can approximate sparse distributions with a sufficiently high CV (we call such patterns ‘sparse-like’).
To summarize, we are interested in how the maximal number of patterns $P$ that can be stored as fixed points, and the fraction of these fixed points that are dynamically stable, depend on parameters $\sigma$, $n$, $\theta$, and CV (as well as $f$, the sparseness of patterns, Details). For obtaining this maximum, we optimize $\mathbf{W}$ following Eq. 4.
@@
A replica analysis of the solution for the optimization problem Eq. 4 follows the Gardner approach [11, 12, 18], without assuming specific connection weights as in the Hopfield model [2, 5]. In the Appendix, we derive a mean-field theory for the capacity to store graded, random memory patterns as fixed points of Eq. 1, without assumptions on the weights. The analysis applies to either dense patterns with a strictly monotonic activation function or to sparse patterns with a rectified monotonic activation function. It generalizes previous work in which a rectified-linear activation function and normalized connection weights were assumed [12]. As usual, our analysis assumes $P,N\to\infty$ with a fixed memory load $\alpha=P/N$ and uses the replica technique, assuming the replica symmetry ansatz [19].
In line with previous results [11], our theory predicts that there is a critical capacity $\alpha_{\mathrm{C}}$, such that as $\alpha$ approaches $\alpha_{\mathrm{C}}$, there is no longer a solution to Eq. 4. We find that $\alpha_{\mathrm{C}}$ depends only on pattern sparseness (Details, Fig. S1g) and is independent of other pattern statistics (CV; Fig. 1f), neural thresholds ($\boldsymbol{\theta}$; Eq. 2, Fig. S3a), and activation function details (smoothness $\sigma$ and exponent $n$; Eq. 3, Fig. 1a), thus extending results from neurons with binary [11, 4] or rectified-linear activation functions (corresponding to $\sigma=0$, $n=1$) to our broader class of activation functions ($\sigma\geq 0$, $n>0$) [12]. Results for the sparse case are presented in Fig. S1 (Details). In the dense case, the critical capacity is $\alpha_{\mathrm{C}}=1$ (again, independent from any of the parameters), and theory predicts the following moments for each element of the connection weight matrix:
@@
For a subcritical load, $\alpha=P/N<\alpha_{\mathrm{C}}$, theory can be compared with numerical experiments by directly solving Eq. 4 using off-the-shelf optimizers [20]. The theory provides a good match for both the mean ($N\,\left\langle W\right\rangle$ from Eq. 7; Fig. 1b,g, Fig. S1c, and Fig. S3b) and the L2-norm of a row of the connection weight matrix ($\sqrt{N\,\left\langle W^{2}\right\rangle}$ from Eq. 6; Fig. 1c,h, Fig. S1d, and Fig. S3c). As $\alpha$ approaches $\alpha_{\mathrm{C}}$, the theory predicts the L2-norm of weights to diverge, explaining why there is no solution to Eq. 4 in this regime (Fig. 1e,j, Fig. S1e,f, and Fig. S3e).
To test if the resulting dynamics might coincide with those assumed by the ‘Hebbian’ approach, we measured the weight matrix asymmetry and the deviation from normality of the dynamics around the fixed points (Details). Both are non-negligible and increase with the problem’s load $\alpha$ and pattern variation (Fig. 1d,i, and Fig. S3d), indicating the resulting dynamics are distinct from those of energy-based models.
Importantly, investigating the emergent stability of patterns in the subcritical regime reveals a second phase transition: all patterns tend to be stable up to a new critical value $\alpha_{\mathrm{S}}<\alpha_{\mathrm{C}}$, above which all patterns tend to be unstable (Fig. 1e,j, Fig. S2a,b, and Fig. S3e). In the following, we investigate this phenomenon more closely.
@@
where $\mathbf{R},\mathbf{V}\in\mathbb{\mathbb{R}}^{N\times P}$ are the stored memory patterns $R_{i\mu}=r_{i}^{\mu}$ and $V_{i\mu}=g^{-1}(R_{i\mu})+\theta$. A slightly more complicated expression is available when avoiding self-couplings, Eq. 19, but we find that we can neglect this difference and use Eq. 8 in the derivation of stability. This generalizes earlier results derived for storing binary patterns, in which case weights defined through the patterns’ pseudo-inverse were already proposed in the ‘Hebbian’ approach [21, 3]. However, those results were limited to a linear activation function $\mathbf{V}=\mathbf{R}$, rendering $\mathbf{W}$ symmetric (and thus allowing the construction of an energy function). In the general case of a non-linear activation function, $\mathbf{W}^{*}$ obtained from Eq. 8 is not symmetric and thus falls outside the scope of the ‘Hebbian’ approach.
Figure 2: Fixed-point stability. (a) Example spectrum - overlaid eigenvalues of $P$ Jacobians $\mathbf{J_{\mu}}$ of a single problem (parameters indicated by text). Annotations of the theoretical values $\lambda_{\mathrm{bulk}}$ (blue line), $\lambda_{\mathrm{ave}}$ (green line), and $\lambda_{\mathrm{mem}}$ (orange line). The corresponding empirical eigenvalues are marked with matching colors: largest eigenvalue of the bulk (blue dot), the eigenvalue associated with the Jacobian average (green dot), and the eigenvalue associate with the memory pattern (orange dot). Inset shows a near-linear relation between $r^{\mu}_{i}$ (x-axis) and $\phi(r^{\mu}_{i})$ (y-axis), a prerequisite for an outlier $\lambda_{\mathrm{mem}}$. (b) Same as a, for six parameter sets; columns differ by the maximal theoretical value. (c-e) Probability densities at different CV values (color-coded) of the difference between the empirical and predicted value of $\lambda_{\mathrm{bulk}}$ (c), $\lambda_{\mathrm{ave}}$ (d), and $\lambda_{\mathrm{mem}}$ (e), calculated over different choice of the parameters discussed in Figs. 1 and S3, in cases where $\left|\lambda_{\mathrm{bulk}}\right|<10$ (c), $\left|\lambda_{\mathrm{ave}}\right|<10$ (d), $\tau_{\mathrm{mem}}>0.95$ (e).
To characterize $\alpha_{\mathrm{S}}$, the critical load for the stability of $\mathbf{J}_{\mu}$ with $\mathbf{W}=\mathbf{W}^{*}$, we note that from Eqs. 5 and 8:
$\mathbf{J}_{\mu}^{*}=-\mathbf{I}+\underbrace{\mathbf{g^{\prime}}\!\left(\mathbf{g}^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)\circ\mathbf{V}}_{\displaystyle\bar{\mathbf{V}}_{\mu}}\,\mathbf{R}^{\dagger}$ (9)
and build on our previous analysis of the eigenvalue spectrum of random matrices $\mathbf{M}=-\mathbf{I}+\mathbf{X}\,\mathbf{Y}^{\dagger}$ when pairs of the corresponding entries of $\mathbf{X}$ and $\mathbf{Y}$ are (jointly) independent and identically distributed (i.i.d.), correlated Gaussian, and $\alpha$ simply denotes the dimension-ratio (width/height) of $\mathbf{X}$ and $\mathbf{Y}$ (a generalization of $\alpha$ being the load in our case) [22]. This analysis provides an exact result for the support of the eigenvalue spectrum (Fig. S5) and the largest real eigenvalue $\lambda_{\mathrm{bulk}}$ of $\mathbf{M}$ (Details, Eq. 21). Interestingly, we find that, just as the Jacobians of the original network, $\mathbf{J}_{\mu}^{*}$, the Jacobians defined by $\mathbf{M}$ (for any choice of $\mathbf{X}$ and $\mathbf{Y}$) also undergo a phase transition: they are either all stable (for $\alpha<\alpha_{\mathrm{S}}$) or all unstable (for $\alpha>\alpha_{\mathrm{S}}$) (Details, Eq. 22). To apply the general theory of the eigenvalue spectrum of $\mathbf{M}$ to the eigenvalue spectrum of $\mathbf{J}_{\mu}^{*}$ (Eq. 9), we choose $\mathbf{X}$ and $\mathbf{Y}$ such that the variances and covariance of $X_{i\mu}$ and $Y_{i\mu}$ are respectively $c_{\mathrm{ff}}=\langle\delta f\!\left(r,r^{\prime}\right)^{2}\rangle$, $c_{\mathrm{rr}}=\langle\delta r^{2}\rangle$, and $c_{\mathrm{rf}}=\left\langle\delta r\,\delta f\!\left(r,r^{\prime}\right)\right\rangle$, where $f\!\left(r,r^{\prime}\right)=g^{\prime}\!\left(g^{-1}\!\left(r^{\prime}\right)\right)\,\left(g^{-1}\!\left(r\right)+\theta\right)$ for independent variables $r$, $r^{\prime}$. With these substitutions,
$\lambda_{\mathrm{bulk}}$ $=-1+\frac{c_{\mathrm{rf}}}{c_{\mathrm{rr}}}+\frac{1}{c_{\mathrm{rr}}}\sqrt{\frac{\alpha}{1-\alpha}\,\left(c_{\mathrm{rr}}\,c_{\mathrm{ff}}-c^{2}_{\mathrm{rf}}\right)}$ (10)
Solving for the largest value of $\alpha$ for which $\lambda_{\mathrm{bulk}}<0$, the resulting critical load for stability from this analysis is:
@@
The foregoing stability analysis was based on the zero-crossing of the ‘rightmost point’ of the bulk of the eigenvalue spectrum of $\mathbf{M}$, $\lambda_{\mathrm{bulk}}$. However, we note that this eigenvalue spectrum can only be an approximation to that of $\mathbf{J}_{\mu}^{*}$ (Fig. 2a), since the entries of $\mathbf{J}_{\mu}^{*}$ are neither i.i.d. (due to the $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r^{\mu}}\right)\right)$ term in $\bar{\mathbf{V}}_{\mu}$, which couples different rows), nor Gaussian (since $\mathbf{V}$ is a deterministic function of $\mathbf{R}$, rather than being a jointly distributed random variable). Despite these facts, we empirically find that the eigenvalue spectrum of $\mathbf{M}$ often provides an acceptable approximation to at least the bulk of the eigenvalue spectrum of $\mathbf{J}_{\mu}^{*}$, so that $\lambda_{\mathrm{bulk}}$ often predicts stability well (Fig. 2a,c and Fig. 2b, left column). Nevertheless, there are also cases when $\lambda_{\mathrm{bulk}}$ alone is an imperfect predictor of stability (Fig. 2b, middle and right columns). Thus, we analyze two specific outlier eigenvalues that might interfere with stability.
The analysis of each of the two outlier eigenvalues is based on assuming the existence of a specific eigenvector of $\mathbf{J}_{\mu}^{*}$ and computing its corresponding eigenvalue, which may be an outlier. In the first case, we make the assumption that the uniform vector, $\boldsymbol{1}$, is an eigenvector. If such an eigenvector existed, the corresponding eigenvalue, $\lambda_{\mathrm{ave}}^{\mu}$, would be the (average) row-sum of $\mathbf{J}_{\mu}^{*}$, $\bar{J}$. Thus, using replica theory (Eq. 7) to compute $N\,\langle W\rangle$, the expectation of this outlier eigenvalue is:
@@
$\lambda_{\mathrm{ave}}=\left\langle\lambda_{\mathrm{ave}}^{\mu}\right\rangle=\left\langle\bar{J}\right\rangle=-1+\left\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\right\rangle\,\frac{\theta+\left\langle g^{-1}\!\left(r\right)\right\rangle}{\langle r\rangle}$ (12)
where by taking the expectation over $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)$, we have again ignored its memory pattern-dependence. Even though the uniform vector is not an eigenvector of $\mathbf{J}_{\mu}^{*}$ in general, when we empirically identify the closest eigenvector (Details), we find that the associated eigenvalue is often close to $\lambda_{\mathrm{ave}}$ (Fig. 2a,d), and can determine the dynamical stability of $\mathbf{J}_{\mu}^{*}$ (Fig. 2b, middle column).
In the second case, we assume that there exists some constant $c$ for which $\mathbf{r}^{\mu}+c\,\boldsymbol{1}$ is an eigenvector of $\mathbf{J}_{\mu}^{*}$. In this case, the following equation must hold (Details):
$\boldsymbol{\phi}\!\left(\mathbf{r}^{\mu}\right)=\left(\lambda_{\mathrm{mem}}^{\mu}+1\right)\,\mathbf{r}^{\mu}+c\,\left(\lambda_{\mathrm{mem}}^{\mu}-\lambda_{\mathrm{ave}}\right)\,\boldsymbol{1}$ (13)
where $\phi_{i}\!\left(\mathbf{\mathbf{r}}\right)=\phi\!\left(r_{i}\right)=f\!\left(r_{i},r_{i}\right)$. In practice, Eq. 13 may not hold exactly, so more generally we can characterize the (approximately) linear relationship between $\mathbf{r}^{\mu}$ and $\boldsymbol{\phi}\!\left(\mathbf{r}^{\mu}\right)$ by the expectation of its correlation coefficient, $\tau_{\mathrm{mem}}^{\mu}$, and slope, $\lambda_{\mathrm{mem}}^{\mu}$:
$\tau_{\mathrm{mem}}$ $=\left\langle\tau_{\mathrm{mem}}^{\mu}\right\rangle=c_{\mathrm{r\phi}}/\sqrt{c_{\mathrm{rr}}\,c_{\mathrm{\phi\phi}}}$ (14)
$\lambda_{\mathrm{mem}}$ $=\left\langle\lambda_{\mathrm{mem}}^{\mu}\right\rangle=-1+c_{\mathrm{r\phi}}/c_{\mathrm{rr}}$ (15)